OpenStudy (anonymous):

Find the second derivative of the function. s=(t^6+3t+3)/ t^2

OpenStudy (mertsj):

It might be easier to write it like this: \[s=t^4+3t ^{-1}+3t ^{-2}\]

OpenStudy (sasogeek):

@iplayffxiv can you continue from the help @Mertsj has provided?

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

i dont understand that

OpenStudy (sasogeek):

He simplified the equation first. now it will be easier to find the derivatives :) still lost? here's a simple way to find a derivative: if \(\large y=ax^n \) the derivative \(\large dy \), is equal to \(\large dy=nax^{n-1} \)

OpenStudy (sasogeek):

\(\large a \) is a constant, if \(\large a \) is 1, it is not written

OpenStudy (anonymous):

does not make sense to me. it is multiple choice. can you choose the right answer?

OpenStudy (sasogeek):

first derivative: \(\large s=4t^3-3t^{-2}-6t^{-3} \) apply the same trick to see if you get ur answer.... take ur time, shouldn't be that difficult :)

OpenStudy (anonymous):

A) d2s/dt2 = 4t^3 -3/t2 -6/t3 B) d2s/dt2 = 12t^4 +6/t +18/t2 C) d2s/dt2 = 4t^2 -3/t3 -6/t4 D) d2s/dt2 = 12t^2 +6/t^3 +18/t4

OpenStudy (anonymous):

you see?

OpenStudy (sasogeek):

find the derivative of \(\large 4t^3 \), your answer will be obvious :) remember, the derivative of \(\large ax^n =nax^{n-1} \)

OpenStudy (anonymous):

i dont know what to do

OpenStudy (sasogeek):

do well to read on how to find derivatives, but the answer you're looking for is the last option

OpenStudy (anonymous):

someone give me answer

OpenStudy (mertsj):

Do you want to learn to do this or do you just want an answer?

OpenStudy (anonymous):

i would like to have the answer

OpenStudy (mertsj):

Then we are all leaving because OS is not an answering service. Goodbye and good luck. All the time that has been spent on this problem was ample for you to learn to do i had you so chosen.

OpenStudy (loser66):

good luck bye!!!